# Keyword-driven Retrieval-Augmented Large Language Models for Cold-start User Recommendations

### v1

@@ 3.2.2. Message Passing on Graph (MPG)
Utilizing CBR alone is not sufficient to provide satisfactory embeddings for Collaborative Filtering. The primary deficiency lies in the embedding function’s failure to encode crucial collaborative signals latent in keyword-restaurant and keyword-user interactions. Drawing inspiration from LightGCN (He et al., 2020), which leverages structured information, we propose a model based on a tripartite graph structure (user - keyword - restaurant). Specifically, we construct the graph using training data, with edges established for (user - keyword) and (keyword - restaurant) interactions when a user uses a keyword to review and a restaurant’s review contains that keyword. Information is passed through the edges in the graph, and node information could be generated as follows:
@@
Figure 3. Illustrate graph construction with connections between users, keywords, and restaurants. Green nodes represent keywords from the training data, while dashed lines indicate connections established via nearest neighbors for keywords not in the data. Edge brightness reflects the importance score.
For a restaurant, the message is generated by multiplying the information of the keyword (source node) by its edge weight, then summing up the messages to update as follows:
@@
where $e^{w}$ denotes feature of keyword nodes $w$; $a^{w}_{r}$ denote the connection between node $w$ and node $r$. Specifically, the edge between a keyword node and a restaurant can be represented by a value indicating the importance of a keyword to a restaurant, denoted as $\mathcal{A}_{r}^{w}$. We introduce a scheme named TF-IRF to measure the importance score of a keyword to a restaurant, which is similar to the TF-IDF idea. The score assigned to the edge established by a restaurant $r$ and a keyword $w$ is denoted as $s^{w}{r}$ and defined as:
@@
in which:
  • $tf_{r}^{w}=\frac{f_{r}^{w}}{q_{w}}$, where $f_{r}^{w}$ represents the number of times the term $w$ appears in the set of keywords for restaurant $r$, and $q_{w}$ indicates the total number of times the term $w$ is used in the entire training data.
  • $irf^{w}=\log\left(\frac{|\mathcal{R}|}{f^{w}}\right)$, where $f^{w}$ denotes the number of restaurants containing the term $w$.
The scores of all restaurants $\mathcal{S}$ can be rewritten in matrix form as:
@@
where $\mathcal{S}\in\mathbb{R}^{1\times n},\mathcal{A}_{r}^{w}\in\mathbb{R}^{w\times r}$, and $\mathcal{M}_{w}\in\mathbb{R}^{1\times w}$ denotes the occurence matrix represent the connection between cold-start user and their selected keywords. Then, the top-k possible restaurants are provided by:
@@ 3.3. LLM-based Candidates Re-ranking
Output: Must include 15 restaurants in the candidate restaurant set, not keyword. Desired format is string: restaurant_id_1, restaurant_id_2, … We utilize LLMs to re-rank candidates for each user, followed by natural language instructions. Additionally, leveraging their reasoning and generation abilities, we incorporate information about user and restaurant candidates into the instructions to make LLMs aware of user preferences. Like the retrieval model, user and restaurant information is represented by keywords. We also include sentences in Template $\mathcal{T}$ to trigger the recommender abilities of LLMs (Hou et al., 2024) and to describe the task instructions to the model:
Following, we propose a general restaurant recommendation prompt framework using keywords, consisting of (1) user keywords; (2) candidate restaurant sets; and (3) restaurant keyword sets. This prompt framework is denoted as [$\mathcal{H}_{z}$] and [$\mathcal{H}_{f}$] represent for zero-shot and few-shot pattern, respectively.
@@ 3.3.1. Zero-shot $\mathcal{H}_{z}$
Since large language models trained on vast, information-rich datasets, zero-shot prompts require meaningful keywords for LLMs to process information and deliver high-quality re-ranking results. When keywords associated with the user are given, LLMs can grasp and anticipate the user’s interests. Therefore, the significance of keywords yields valuable insights. In the same way, the LLMs also comprehends the characteristics of restaurants, enabling it to recommend that are relevant and appealing to users based solely on the keywords provided about the restaurant.
@@ 3.3.2. Few-shot [$\mathcal{H}_{f}$]
The pattern of few-shot prompts [$\mathcal{H}_{f}$] above comprises examples and the information in [$\mathcal{H}_{z}$] then filled into Template $\mathcal{T}$. The keyword candidates are ordered in our prompts to preserve the order from retrieval models, which is based on TF-IRF scores.
Examples are generated from users in the training set. For each example, candidates are selected based on the number of overlapping keywords to a chosen user. The ranked list for each user is obtained by sorting the rating of the candidates.
@@ 4. Experiments
In this section, we conduct experiments on a real-world dataset from various cities to evaluate the performance of our KALM4REC framework. Specifically, we assess the candidate retrieval module and the LLMs’ capabilities in re-ranking restaurant candidates for cold-start user recommendations. Additionally, we report an extensive set of experiments to investigate the impact of various factors on the quality of the LLMs’ re-ranking.
@@ 4.1. Experimental Settings
Dataset. The experiments utilize the Yelp.com dataset (Vu et al., 2020), featuring over 67,000 restaurant reviews across three English-speaking cities. The dataset is partitioned into training and testing sets. Since we deal with the problem of cold-start user recommendation, we split the dataset such that users in the test set do not have any reviews in the train set. Dataset statistics are provided in Table 1.
Table 1. Data statistic of the Yelp dataset which contains reviews from three English-speaking cities. We split the train/test sets such that no users in the test set has review in the train set (cold-start user setting).
City # Restaurants # Reviews # Users
@@
City Conventional Ours Improve
@@
Figure 4. Examine other impacts to KALM4REC . Left figure indicate effectiveness of keyword score, ”w” refers result for TF-IRF score mentioned in 6 while ”1” refers all scores are 1. Right figure indicate performance of different BERT model to encode the keyword for MPG to find the closest keyword in case of there is no matching keyword in training data. ”1” refers to all-MiniLM-L6-v2 while ”0” indicates whaleloops/phrase-bert. Both results are report in Precision@20 using MPG.
@@
Retrieval Baselines. To evaluate our retrieval module, we compared it with the following retrieval baseline methods:
  • Jaccard similarity-based method measures the similarity between cold-start user and restaurants based on the overlapping keywords. The restaurants candidates are ranked based on Jaccard similarity scores.
  • Matrix Factorization (MF) is a classic collaborative filtering method that decomposes the user-item interaction matrix and use the BPR (Rendle et al., 2009) loss to optimized. The optimized function is described in Equation 3. After training, the model learns the relationships between the users in the training set and the restaurants. For cold-start user recommendation, we first retrieve similar users from the training set using Jaccard similarity based on the input keywords. Then, we rank the restaurants based on the average ranking scores of these similar users.
  • Multi-Variational Autoencoder (MVAE) is an item-based Collaborative Filtering method based on the variational autoencoder. It assumes that the data is generated from a multinomial distribution and uses variational inference for parameter estimation. We utilize the code provided by the authors [^8] https://github.com/younggyoseo/vae-cf-pytorch. For generating candidates in cold-start user recommendations, we employ the same strategy as with MF: the candidate scores are the average scores from similar training users, determined using Jaccard similarity based on input keywords.
Table 3. Performance in comparison with different retrieval models with Gemini. All ALMs obtain remarkable improvement against it backbone in retrieval module that indicate the effectiveness of our framework. P, R, N stands for Precision, Recall, and NDCG, respectively.
City Methods P@1 R@1 P@3 R@3 N@3 P@5 R@5 N@5
Edinburgh CBR 29.15 5.05 19.13 10.23 32.81 15.07 13.19 32.87
@@
MPG-LLM 55.15 11.19 36.48 20.48 55.40 28.53 26.89 54.03
@@ 4.2. Evaluating Candidates Retrieval
The goal of candidates retrieval is not only to achieve a high ranking (precision) but also to retrieve as many correct candidates as possible (recall). This ensures that the re-ranking module has a comprehensive set of candidates to work with in its input. Therefore, we evaluate the retrieval methods using precision and recall of the top 20 candidates, i.e., Precision@20 and Recall@20. The results are reported in Table 2. Column “Improve” indicates the improvement of our best retrieval model (MPG) against the best conventional approach. Table 2 shows that both MF and MVAE outperform Jaccard similarity-based method. However, these methods face challenges because they are trained on user-restaurant interactions rather than keywords, making it difficult to accurately capture the intentions of cold-start users as expressed through input keywords. In contrast, approaches leveraging keywords exhibit notable enhancements. CBR performs better than all the baselines method as it utilized keywords to compute user and restaurant representations. The best performance is achieved using MPG, which harnesses the graph structure between keywords and restaurants. MPG demonstrates consistent improvement across three cities, enhancing Recall@20 by at least 21.38% and Precision@20 by at least 4.75%. Furthermore, we examine the role of TF-IRF in constructing edges( Equation 6) and pretrained BERT model for the keyword vectorization. As shown in Figure 4, MPG with TF-IRF and ”all-MiniLM-L6-v2” outperforms other configurations.It is noteworthy that in our framework, achieving high recall is crucial, as these candidates will be re-ranked in the next phase.
@@ 4.3. Investigating the Re-ranking Capability of LLMs for Cold-start User Recommendations
  • RQ1. Can LLMs re-rank candidates to improve the cold-start user recommendation task?
  • RQ2. Is employing keywords more effective than using reviews in prompting?
  • RQ3. How do prompt design strategies impact the efficiency of LLMs?
  • RQ4. The impact of keywords on the re-ranking performance.
  • RQ5. Do LLMs suffer from biases while ranking?
RQ1. Can LLMs re-rank candidates to improve the cold-start user recommendation task? We investigate the capability of using LLMs to re-rank candidates in cold-start user recommendation. Table 3 shows the comparison between retrieval only (CBR and MPG) and the full KALM4REC framework (i.e., retrieval and re-ranking using LLM – Gemini Pro). The results show the consistent improvements when using LLM-based candidates re-ranking for both CBR and MPG, for all the 3 cities regarding the evaluation metrics. Additionally, there is a significant improvement compared to sole retrieval modules, indicating the extensibility of our framework with alternative retrieval modules in the future.
Table 4. Performance of KALM4REC using different LLMs for re-ranking with MPG as the retrieval method. The results demonstrate that all LLMs enhance the precision and recall of the top recommended item, with Gemini consistently delivering the best overall performance. P, R, N stands for Precision, Recall, and NDCG metrics, respectively.
@@
Figure 5. Comparing the recommendation performance between the candidates retrieval module (MPG) and KALM4REC with the use of reviews or keywords. Results show that prompting using keywords obtain consistent improvements over MPG, whereas prompting using reviews only shows improvement at P@1 and R@1. Experiment is conducted using Gemini Pro.
Figure 6. Total number of prompt tokens when using reviews versus keywords to prompt Gemini Pro (in millions).
We also examine the performance of KALM4REC with different LLMs. Table 4 compares the performance of various LLM models –Gemini, GPT-3.5, Mistral 8B, and Llama3 8B– across multiple cities using precision at 1 (P@1), recall at 1 (R@1), precision at 3 (P@3), recall at 3 (R@3), and NDCG at 3 (N@3). We use MPG as the retrieval method. The results demonstrate that all LLMs enhance the precision and recall of the top recommended item (P@1 and R@1). Across all cities, Gemini consistently achieves the highest scores in almost every metric, demonstrating its robustness in identifying and ranking relevant restaurant candidates. For instance, Gemini attains the highest P@1 values, indicating superior accuracy in recommending the top restaurant. Additionally, Gemini’s strong performance in N@3 reflects its ability to rank relevant candidates effectively. GPT-3.5 and Mistral 8B models also demonstrate improvements over MPG. Both models show competitive results in P@3 and R@3, indicating their effectiveness in providing multiple relevant recommendations. GPT-3.5, for example, often comes close to Gemini’s performance, particularly in N@3, suggesting its capability to enhance the ranking quality significantly compared to MPG. The Mistral 8B model, while not always leading, maintains consistent and reliable performance across all metrics. Llama 3-8B, though slightly behind the other LLMs, still shows improvement over MPG in several metrics. While it may not be as effective in capturing the top preferences of users as the leading models, it provides a consistent alternative that surpasses the baseline.
RQ2. Is employing keywords more effective than using reviews in prompting? To convey user preferences and restaurant characteristics, the straightforward approach is to use the full content of user reviews. However, due to the context length limitations in LLM prompts, KALM4REC uses keywords instead. We compare the effectiveness of LLMs when using reviews versus keywords for the re-ranking task. For each restaurant candidate in a prompt, we describe it using a representative review, which is selected based on its distance to the review centroid. The centroid embedding is calculated as the average of the embeddings of all the restaurant’s reviews. We use Gemini Pro for this evaluation. The comparison results across 3 cities are illustrated in Figure 5 in Precision@k and Recall@k with k = 1, 3. From Figure 5, we find that rankings by our KALM4REC consistently show remarkable improvements across all cities in P@1 and R@1. However, we observe KALM4REC with review gets lower number in P@3 and R@3 compared to the retrieval model for Singapore. Whereas, using keywords offers improvements across all the metrics. This improvement is likely attributed to the detailed information provided by keywords, which directly correspond to the input keywords. This enables the LLM to better understand the relevance between the input user and the candidate restaurants.
Another benefit of using keywords is the length-efficiency when prompting LLMs. Figure 6 depicts the total number of prompt tokens used when utilizing reviews versus keywords to prompt Gemini Pro. Using reviews required approximately 800,000 tokens for Singapore, 950,000 for Edinburgh, and 3,190,000 for London. In contrast, using keywords required only about 170,000, 340,000, and 718,000 tokens for the respective cities. In total, prompts using reviews are four times longer than those using keywords. Keywords not only outperform reviews in effectiveness but are also much more efficient in terms of prompt length and associated costs.
RQ3. How do prompt design strategies impact the efficiency of LLMs? We compare the performance of KALM4REC with zero-shot [$\mathcal{H}_{z}$] and few-shot [$\mathcal{H}_{f}$] strategies. In this experiment, we set number of keywords for user ranging from 3 to 5 while number of keywords for restaurant range from 3 to 20. We report the best setting for each strategy in Figure 7. The experiments is conducted with Gemini Pro. We observe significant performance improvements across all strategies of zero-shot and few-shot when compared to MPG. The number of shots (1 shot, 2 shots, 3 shots) indicates the number of examples introduced to LLMs. Overall, increasing the number of examples enhances LLM performance, with the 3-shot approach mostly achieve the best results. This outcome is reasonable as LLMs can better capture user intent with more examples.
RQ4. The impact of keywords on the re-ranking performance. We investigate the impact of keywords on the re-ranking performance regarding the number of keywords and their orders. Specifically, we compare the re-ranking results when varying the number of keywords used to describe restaurants and users. We also examine whether the order of keywords affect the performance of KALM4REC. Figure 8 shows the Recall@1 of zero-shot and one-shot for Singapore using Gemini Pro. We observed that one-shot requires fewer keywords yet still achieves relatively good performance. Figure 9 shows the evaluation performed on Edinburgh and Singapore using Gemini Pro when we change number of keywords and whether to shuffle the order of the keywords in prompts. The results show that the number of keywords used for restaurants influences LLM performance. Specifically, in Edinburgh, LLM performance increases with the number of keywords (with order). The visualizations also emphasize the crucial role of keyword order in the prompt. Performance notably declines when keywords are not ordered according to our ranking, indicating that LLMs better comprehend user preferences when keywords are sorted by their importance score. This effect is particularly pronounced in Edinburgh, where performance without the ordered keywords is even lower than that of MPG.
@@
Figure 7. The impact of different ranking strategies on performance across various cities. The red dashed line represents results from the retrieval model only. Experiement is conducted on Gemini.
@@
Figure 8. Comparison the number of keywords used to describe restaurant candidates and users in zero-shot and one-shot scenarios in Recall@1, we find that one-shot requires fewer keywords yet still achieves relatively good performance. Experiment is conducted using Gemini Pro on Singapore data.
(a) Edinburgh
(b) Singapore
Figure 9. Performance comparison in terms of the number of keywords and their order of representation for each restaurant within the zero-shot design. Experiment is conduct with Gemini for Edinburgh.
RQ5. Do LLMs suffer from biases while ranking?
We adopt the insights from (Hou et al., 2024) that the order of candidates affects the ranking results of LLMs. As the result, we make an experiment where the order of restaurant candidates is shuffled. The experiment is conducted with Gemini Pro and depicted in Figure 10. The shown results indicate that Gemini are generally sensitive to the order of examples in the prompt. Especially, the ranking performance by shuffle drops significantly for Singapore and London, lower than performance of retrieval models. This suggests that LLMs perform well when integrated with efficient retrieval models, as demonstrated in our previous experiment (Table 4), highlighting the crucial role of our proposed retrieval model.
Figure 10. Impact of candidate order in prompt to our KALM4REC. Experiment is conducted using Gemini Pro.

### v2

@@ 3.2.2. Message Passing on Graph (MPG)
Utilizing CBR alone is not sufficient to provide satisfactory embeddings for Collaborative Filtering. The primary deficiency lies in the embedding function’s failure to encode crucial collaborative signals latent in keyword-item and keyword-user interactions. Drawing inspiration from LightGCN (He et al., 2020), which leverages structured information, we propose a model based on a tripartite graph structure (user - keyword - item). Specifically, we construct the graph using training data, with edges established for (user - keyword) and (keyword - item) interactions when a user uses a keyword to review and a item’s review contains that keyword. Information is passed through the edges in the graph, and node information could be generated as follows:
@@
Figure 3. Graph construction with connections between users, keywords, and items. Green nodes represent keywords from the training data, while dashed lines indicate connections established via nearest neighbors for keywords, not in the data. Edge brightness reflects the importance score.
For an item, the message is generated by multiplying the information of the keyword (source node) by its edge weight, then summing up the messages to update as follows:
@@
where $e^{w}$ denotes feature of keyword nodes $w$; $a^{w}_{r}$ denote the connection between node $w$ and node $r$. Specifically, the edge between a keyword node and an item can be represented by a value indicating the importance of a keyword to the item, denoted as $\mathcal{A}_{r}^{w}$. We introduce a scheme named TF-IRF to measure the importance score of a keyword to an item, which is similar to the TF-IDF idea. The score assigned to the edge established by an item $r$ and a keyword $w$ is denoted as $a^{w}{r}$ and defined as:
@@
with $tf_{r}^{w}=\frac{f_{r}^{w}}{q_{w}}$, where $f_{r}^{w}$ represents the number of times the term $w$ appears in the set of keywords for item $r$, and $q_{w}$ indicates the total number of times the term $w$ is used in the entire training data; $irf^{w}=\log\left(\frac{|\mathcal{R}|}{f^{w}}\right)$, where $f^{w}$ denotes the number of items containing the term $w$. The scores of all items $\mathcal{S}$ can be rewritten in matrix form as:
@@
where $\mathcal{S}\in\mathbb{R}^{1\times r},\mathcal{A}_{r}^{w}\in\mathbb{R}^{w\times r}$, and $\mathcal{M}_{w}\in\mathbb{R}^{1\times w}$ denotes the occurence matrix represents the connection between cold-start user and their selected keywords. Then, the top-k possible items are provided by:
@@ 3.3. LLM-based Candidates Re-ranking
Output: Must include 15 restaurants in the candidate restaurant set, not keyword. Desired format is string: restaurant_id_1, restaurant_id_2, …
Following, we propose a general item recommendation prompt framework using keywords, consisting of (1) user keywords; (2) candidate sets; and (3) item keyword sets. This prompt framework is denoted as [$\mathcal{H}_{z}$] and [$\mathcal{H}_{f}$] represent for zero-shot and few-shot pattern, respectively.
@@ 3.3.2. Few-shot [$\mathcal{H}_{f}$]
Similar to human reasoning, LLMs benefit from good examples to better understand human intentions and criteria for desired answers. Therefore, few-shot learning often yields better performance than zero-shot. The pattern of few-shot prompts [$\mathcal{H}_{f}$] above comprises examples and the information in [$\mathcal{H}_{z}$] then filled into Template $\mathcal{T}$. The keywords of candidates are ordered in our prompts to preserve the order from retrieval models, which is based on TF-IRF scores.
@@ 4. Experiments
In this section, we conduct experiments on a real-world dataset from various cities to evaluate the performance of our KALM4REC framework. Specifically, we assess the candidate retrieval module and the LLMs’ capabilities in re-ranking candidates for cold-start user recommendations. Additionally, we report an extensive set of experiments to investigate the impact of various factors on the quality of the LLMs’ re-ranking.
@@ 4.1. Experimental Settings
Dataset. The experiments utilize two datasets: the Yelp.com dataset (Vu et al., 2020) and the TripAdvisor dataset (Li et al., 2013). The Yelp.com dataset features over 67,000 restaurant reviews across three English-speaking cities. The TripAdvisor dataset consists of 878,561 reviews from 4333 hotels crawled from TripAdvisor.com across 17 cities. The dataset is partitioned into training and testing sets. Since we deal with the problem of cold-start user recommendation, we split the dataset so that users in the test set do not have any reviews in the train set. Dataset statistics are provided in Table 1.
Table 1. Statistic of the datasets. In the Yelp dataset, ”Items” refers to restaurants, while in the TripAdvisor dataset, ”Items” refers to hotels.”
Dataset City # Items # Reviews # Users
@@
Dataset Conventional Our Retrievals Improve
@@
[…]
Figure 4. The left figure indicates the effectiveness of the keyword score, ”weighted” refers result for TF-IRF score mentioned in Equation 6 while ”all_ones” refers to assigning all edge’s scores to 1. The right figure indicates the performance of different pretrained BERT models to vectorized the keywords. ”MiniLM” refers to all-MiniLM-L6-v2 while ”whaleloops” indicates whaleloops/phrase-bert.
@@
Retrieval Baselines. Retrieval Baselines. To evaluate our retrieval module, we compared it with the following retrieval baseline methods:
  • Jaccard similarity-based method measures the similarity between cold-start users and items based on the overlapping keywords. The candidates are ranked based on Jaccard similarity scores.
  • Matrix Factorization (MF) is a classic collaborative filtering method that decomposes the user-item interaction matrix and uses the BPR (Rendle et al., 2009) loss to optimize.
  • Multi-Variational Autoencoder (MVAE) is an item-based Collaborative Filtering method based on the variational autoencoder. We utilize the code provided by the authors [^8] https://github.com/younggyoseo/vae-cf-pytorch.
For cold-start user recommendation in MF and MVAE, we retrieve similar users from the training set using Jaccard similarity based on the input keywords. Then we estimate the item score for cold-start users based on the average ranking scores of their similar users.
Table 3. Performance in comparison with different retrieval models with Gemini. All LLMs obtain remarkable improvement against their backbone in the retrieval module.
Dataset Methods P@1 R@1 P@3 R@3 N@3 P@5 R@5 N@5
Yelp-Edinburgh CBR 29.15 5.05 19.13 10.23 32.81 15.07 13.19 32.87
@@
MPG-LLM 58.06 11.77 37.09 20.67 56.85 29.18 26.24 54.30
TripAdvisor CBR 21.38 3.35 15.30 7.16 28.92 12.99 10.00 31.23
CBR-LLM 36.27 5.79 20.05 9.42 36.69 14.84 11.49 11.49
MPG 54.60 8.73 41.92 19.72 61.36 33.52 25.98 58.99
MPG-LLM 59.43 9.54 42.15 19.87 62.36 33.58 25.98 59.65
Figure 5. Comparing the performance between the candidates retrieval module (MPG) and KALM4REC with the use of reviews or keywords. Results reported at Precision@1 and Recall@1. The experiment is conducted using Gemini Pro.
@@ 4.2. Evaluating Candidates Retrieval
The goal of candidates retrieval is not only to achieve a high ranking (precision) but also to retrieve as many correct candidates as possible (recall). This ensures that the re-ranking module has a comprehensive set of candidates to work with in its input. Therefore, we evaluate the retrieval methods based on P@20 and R@20, which measure the precision and recall of the top 20 candidates, respectively. The results are presented in Table 2, where the column ”Improve” highlights the performance improvement of our best retrieval model, MPG, compared to the best conventional method. Our findings reveal that both MF and MVAE outperform Jaccard similarity-based method. However, these methods face challenges because they are trained on user-item interactions rather than keywords, making it difficult to accurately capture the intentions of cold-start users as expressed through input keywords. In contrast, approaches leveraging keywords exhibit notable enhancements. CBR performs better than all the baselines method as it utilized keywords to compute user and item representations. The best performance is achieved using MPG, which harnesses the graph structure between keywords and items. MPG demonstrates consistent improvement across three cities, enhancing Recall@20 by at least 21.38% and Precision@20 by at least 4.75%. Furthermore, we examine the role of TF-IRF in constructing edges (Equation 6) and pretrained BERT models for the keyword vectorization. As shown in Figure 4, MPG with TF-IRF and ”all-MiniLM-L6-v2” outperforms other configurations. It is noteworthy that in our framework, achieving high recall is crucial, as these candidates will be re-ranked in the next phase.
@@ 4.3. Investigating the Re-ranking Capability of LLMs for Cold-start User Recommendations
RQ1. Can LLMs re-rank candidates to improve the cold-start user recommendation task? We investigate the capability of using LLMs to re-rank candidates in cold-start user recommendation. Table 3 shows the comparison between retrieval only (CBR and MPG) and the full KALM4REC framework (i.e., retrieval and re-ranking using LLM – Gemini Pro). The results show consistent improvements when using LLM-based candidates re-ranking for both CBR and MPG, for all the datasets regarding the evaluation metrics. Additionally, there is a significant improvement compared to sole retrieval modules, indicating the extensibility of our framework with alternative retrieval modules in the future.
Table 4. Performance of KALM4REC using different LLMs for re-ranking with MPG as the retrieval method. The results demonstrate that all LLMs enhance the precision and recall of the top recommended item, with Gemini consistently delivering the best overall performance.
@@
(a) Number of token
(b) Fewshot performance
@@
Figure 6. Left figure: Total number of prompt tokens when using reviews versus keywords to prompt Gemini Pro (in millions). Right figure: The impact of different ranking strategies on performance across various cities. The dashed line represents results from the retrieval model only.
We also examine the performance of KALM4REC with different LLMs. Table 4 compares the performance of various LLM models –Gemini, GPT-3.5, Mistral 8B, and Llama3 8B– across multiple cities in Yelp dataset using Precision at 1 (P@1), Recall at 1 (R@1), Precision at 3 (P@3), Recall at 3 (R@3), and NDCG at 3 (N@3). We use MPG as the retrieval method. The results demonstrate that all LLMs enhance the precision and recall of the top recommended item. Across all cities, Gemini consistently achieves the highest scores in almost every metric, demonstrating its robustness in identifying and ranking relevant item candidates. For instance, Gemini attains the highest P@1 values, indicating superior accuracy in recommending the top item. Additionally, Gemini’s strong performance in N@3 reflects its ability to rank relevant candidates effectively. GPT-3.5 and Mistral 8B models also demonstrate improvements over MPG. Both models show competitive results in P@3 and R@3, indicating their effectiveness in providing multiple relevant recommendations. GPT-3.5, for example, often comes close to Gemini’s performance, particularly in N@3, suggesting its capability to enhance the ranking quality significantly compared to MPG. The Mistral 8B model, while not always leading, maintains consistent and reliable performance across all metrics. Llama 3-8B, though slightly behind the other LLMs, still shows improvement over MPG in several metrics. While it may not be as effective in capturing the top preferences of users as the leading models, it provides a consistent alternative that surpasses the baseline.
(a) Keyword Order
(b) Candidate Order
(c) Position of Candidate
(d) Popularity Bias
Figure 7. Performance examination of LLMs across various aspects, conducted using Gemini Pro
RQ2. Is employing keywords more effective than using reviews in prompting? To convey user preferences and item characteristics, the straightforward approach is to use the full content of user reviews. However, due to the context length limitations in LLM prompts, KALM4REC uses keywords instead. We compare the effectiveness of LLMs when using reviews versus keywords for the re-ranking task on Gemini Pro. For each item candidate in a prompt, we describe it using a representative review, which is selected based on its distance to the review centroid. The centroid embedding is calculated as the average of the embeddings of all the item’s reviews. The comparison results across 3 cities are illustrated in Figure 5 in Precision@k and Recall@k with k = 1, 3. From Figure 5, we find that rankings by our KALM4REC consistently show remarkable improvements across all cities in P@1 and R@1. However, we observe KALM4REC with review gets a lower number in P@3 and R@3 compared to the retrieval model for Singapore. Whereas, using keywords offers improvements across all the metrics. This improvement is likely attributed to the detailed information provided by keywords, which directly correspond to the input keywords. This enables the LLM to better understand the relevance between the input user and the candidate items.
Another advantage of using keywords is their efficiency in prompting LLMs. Figure 6(a) indicates that prompting with keywords significantly reduces the total number of tokens compared to using reviews on Gemini Pro. This reduction in token length leads to lower associated costs and enhanced efficiency. Consequently, across all datasets, employing keywords not only offer superior effectiveness but also provide a more efficient and cost-effective approach for prompting LLMs.
RQ3. How do prompt design strategies impact the efficiency of LLMs? We compare the performance of KALM4REC with zero-shot [$\mathcal{H}_{z}$] and few-shot [$\mathcal{H}_{f}$] strategies. In this experiment, we set number of keywords for user ranging from 3 to 5 while number of keywords for item range from 3 to 20. We report the best setting with Gemini Pro for each strategy in Figure 6(b). We observe significant performance improvements across all strategies of zero-shot and few-shot when compared to MPG. The number of shots (1 shot, 2 shots, 3 shots) indicates the number of examples introduced to LLMs. Overall, increasing the number of examples enhances LLM performance, with the 3-shot approach mostly achieve the best results. This outcome is reasonable as LLMs can better capture user intent with more examples.
@@
Figure 8. Comparison of the performance of various numbers of keywords used to describe candidates and users at Precision@1. The experiment is conducted using Gemini Pro on Yelp-Singapore data.
RQ4. Do LLMs suffer from biases while ranking? We adopt the insights from (Hou et al., 2024) that suggest several factors can affect the ranking performance of LLMs. To explore potential biases in LLM ranking, specifically on Gemini Pro, we conducted experiments focusing on various aspects. First, we examined how the order of keywords in the prompt affects performance. We shuffled the keywords for each testing user and reported the average performance of re-ranking. As shown in Figure 7a, the performance of re-ranking with shuffled keywords varies, indicating that keyword order influences LLMs. Notably, performance declines when keywords are not sorted by importance, suggesting that LLMs better understand user preferences when keywords are arranged according to their relevance. This effect is especially pronounced in Edinburgh, where performance with unordered keywords is even worse than with MPG.
Additionally, we investigated the impact of candidate order on LLMs. The results, shown in Figure 7b, reveal that Gemini Pro is sensitive to the order of candidates in the prompt. Shuffling the candidate order leads to a significant drop in baseline performance at P@3. However, LLMs’ re-ranking performance shows notable improvement. This suggests that even if the list of candidates is shuffled, LLMs can still recommend suitable items if ground truth items are presented. While the re-ranking performance with shuffled candidates is relatively stable, it remains lower compared to re-ranked MPG. This indicates that integrating efficient retrieval models with LLMs enhances overall performance, as the order of candidates by retrieval positively influences the results.
Next, we assessed LLM performance under the assumption that LLMs can recommend suitable items even if ground truth items are positioned in less favorable locations within the candidate list. We replaced all ground truth items in the candidate list at positions 5, 10, 15, and at the end, ensuring that each list contained no more than four ground truth items to avoid overlap. Results shown in Figure 7c indicate that LLMs are effective at identifying ground truth items even in challenging scenarios, demonstrating relatively stable performance across different settings. However, with the order provided by the retrieval module, KALM4REC achieved the best results highlighting the efficiency of our framework.
Finally, we explore how item popularity affects LLM recommendations. While item popularity cannot be measured directly, we use the normalized interaction count from the training set to represent it. Figure 7d illustrates the relationship between the average popularity score of items and their position in the re-ranked list (e.g., positions 0, 4, 9, 14). We find that highly popular items are often recommended at the top of the re-ranked list, whereas items with lower popularity tend to appear in lower positions. This suggests that popular items have a higher chance of being ranked favorably.
RQ5. The impact of the number of keywords on the re-ranking performance. Figure 8 presents heatmaps depicting the performance of KALM4REC across varying numbers of shots under different keyword conditions for users and items. The x-axis represents the number of keywords used to describe the user, while the y-axis represents the number of keywords used to describe the items. The color intensity reflects the model’s performance. The heatmaps reveal that increasing the number of shots generally enhances accuracy, fewer keywords can still yield relatively good performance. Besides, performance variations are observed depending on the number of keywords for users and items. We also acknowledge that the quality of keywords can affect KALM4REC’s performance. To investigate this, we compared various keyword extractors in Table 5. Our experiments show that SpaCy consistently delivered the best performance in keyword extraction.
Table 5. Comparison performance of our retrieval module with different keyword extractors
Keyword Extractors
spaCy Rake Yake NLTK
Dataset P@20 R@20 P@20 R@20 P@20 R@20 P@20 R@20
[…]